(3.3.15) Based on sample​ data, newborn males have weights with a mean of 3243.4g and a standard deviation of 574.9g. Newborn females have weights with a mean of 3025.6g and a standard deviation of 724.6g. Who has the weight that is more extreme relative to the group from which they​ came: a male who weighs 1700g or a female who weighs 1700​g?'

Since the z score for the male is z= and the z score for the female is z=, the male/female has the weight that is more extreme
1
Expert's answer
2020-11-23T18:24:57-0500

Since the z score for the male is "z = (1700 - 3243.4)\/574.9 = -2.685" and the z score for the female is "z = (1700 - 3025.6)\/724.6 = - 1.829" , the male has the weight that is more extreme